f(x) = (archtanh √3x^2)^3


1
Expert's answer
2021-04-29T17:56:06-0400

"\displaystyle
f(x) = (\mathrm{archtanh}(\sqrt{3x^2}))^3\\

\begin{aligned}
f'(x) &= 3(\mathrm{archtanh}(\sqrt{3}x))^2 \times \sqrt{3} \times \frac{1}{1 - 3x^2} 
\\&= \frac{3\sqrt{3}(\mathrm{arctanh}(\sqrt{3x^2}))^2}{1 - 3x^2} 
\end{aligned}"